OpenStudy (anonymous):

write the expression as a complex number in standard form (8+4i)(2+i)

OpenStudy (anonymous):

We have : \[(8+4i)(2+i)=16+8i+8i+4i^2 \\=16+16i-4 \\=12+16i\]

OpenStudy (anonymous):

thats the final answer

OpenStudy (anonymous):

It should be the last answer as I think ! ;)

OpenStudy (anonymous):

ok thank you